# M-estimation for Gaussian processes with time-inhomogeneous drifts from high-frequency data

### v2

@@ Abstract
We propose a contrast-based estimation method for Gaussian processes with time-inhomogeneous drifts, observed under high-frequency sampling. The process is modeled as the sum of a deterministic drift function and a stationary Gaussian component with a parametric kernel. Our method constructs a local contrast function from adjacent increments, which avoids inversion of large covariance matrices and allows for efficient computation. We prove consistency and asymptotic normality of the resulting estimators under general ergodicity conditions. A distinctive feature of our approach is that the drift estimator attains a nonstandard convergence rate, stemming from the direct Riemann integrability of the drift density. This highlights a fundamental difference from standard estimation regimes. Furthermore, when the local contrast fails to identify all parameters in the covariance kernel, moment-based corrections can be incorporated to recover identifiability. The proposed framework is simple, flexible, and particularly well suited for high-frequency inference with time-inhomogeneous structure.
@@ 2.1 Gaussian processes with time-inhomogeneous drifts
which is not smooth at $t=0$ since it includes $|t|$ in the exponent. However, we can approximate this kernel with a mollifier $\varphi\in C^{\infty}(\mathbb{R})$ such as: ${\rm supp}\,\varphi$ is compact;
@@
$\lim_{n\to\infty}\left[\sum_{i=1}^{\infty}g(t_{i-1})h_{n}-\sum_{i=1}^{n}g(t_{i-1})h_{n}\right]=\lim_{n\to\infty}\int_{nh_{n}}^{\infty}g(s)\,\mathrm{d}s\to 0.$
@@ 3.1 Local-Gauss contrast
Suppose the same conditions as in Theorem 3.1. Suppose further B6 and B8. Then, the estimator $\widehat{\theta}_{n}$ is asymptotically normal under (2.2):
@@
$\Sigma(\theta_{0}):=\begin{pmatrix}\displaystyle-\partial_{t}^{2}K_{\sigma_{0}}(0)\left[2\int_{0}^{\infty}\{\partial_{\xi}\mu_{\xi}(t)\}^{\otimes 2}\,\mathrm{d}t\right]^{-1}&0\\ 0&V_{2}^{-1}(\sigma_{0})V_{1}(\sigma_{0})V_{2}^{-1}(\sigma_{0})\end{pmatrix}$.
@@ 3.2 Moment estimators
$\sqrt{n}\left(\widetilde{\sigma}_{n}-\sigma_{0}\right)\xrightarrow{d}\mathcal{N}\left(\mbox{\boldmath$0$},\ A^{-1}\Gamma^{2}A^{-\top}\right),\quad n\to\infty.$
@@
  • $G(x,y)=(y-x)y^{2}$ combines local variation and level.
@@ 4.1 Drifted Gaussian processes with Gaussian Kernels
For separate estimation of $\alpha$ and $\beta$, we shall consider the method of moment. For example, we can use Lemmas B.5 and B.6 with $f(x,\theta)=x^{2}$ and $G(x,y,\theta)=(y-x)^{2}$ or $(y-x)y^{2}$, respectively: it follows for
@@
$\frac{1}{n}\sum_{i=1}^{n}(Y_{i-1}^{n})^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha;\quad\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}(\Delta_{i}^{n}Y)^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha\beta;\quad\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}\Delta_{i}^{n}Y\cdot(Y_{i-1}^{n})^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}3\alpha^{2}\beta,$
@@ 4.3.2 Discussion of numerical experiments
Future work. Possible directions for improvement are as follows. (1) Random subsample averaging (Jackknife-after-bootstrap type): Generate random subsamples to estimate $\xi$, and use the average of these estimates. This can weaken the correlation between the estimation errors of $\xi$ and $(\alpha,\beta)$ without significantly increasing the variance of $\xi$, thus maintaining stability in finite samples more effectively than simple sample splitting. (2) Orthogonalized estimating functions using lagged covariances: Design the estimating functions for $\alpha$ and $\beta$ so that their gradient with respect to $\xi$ is zero (Neyman orthogonal), thereby ensuring that the estimation error of $\xi$ does not appear as a first-order term. In particular, use covariances at sufficiently large lags, rather than short differences, to extract information on $\alpha$ while attenuating the influence of the drift. (3) One-step stabilization: Starting from stable initial values of Case (I) type (consistent at rate $h_{n}^{-1/2}$ for $\xi$ and $n^{1/2}$ for $\alpha$ and $\beta$), apply a single Newton update to the joint estimating equations. This can correct the first-order dependence between $\xi$ and $(\alpha,\beta)$ and potentially mitigate the bias imbalance observed in finite samples. The goal here is not to improve efficiency, but to achieve better error propagation alignment and numerical stability.
@@ 6 Proofs of main theorems
$D_{n}(\widehat{\theta}_{n}-\theta_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}{\cal N}\left(\mbox{\boldmath$0$},I^{-1}(\theta_{0})J(\theta_{0})I^{-1}(\theta_{0})\right)=\mathcal{N}(0,\Sigma(\theta_{0})).$
@@
$\Phi_{n}(\widetilde{\sigma}_{n})=\mbox{\boldmath$0$}.$
@@
$\sqrt{n}\Phi_{n}(\sigma_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}(\mbox{\boldmath$0$},\Gamma^{2}),\quad n\to\infty,$
@@
$\sqrt{n}\Phi_{n}(\sigma_{0})\xrightarrow{d}\mathcal{N}(\mbox{\boldmath$0$},\Gamma^{2}).$
@@ Appendix A Limit theorems for stationary Gaussian processes
$(Z_{t},Z_{s})^{\top}\sim{\cal N}(\mbox{\boldmath$0$},\Sigma(t,s)),\quad\Sigma(t,s)=\begin{pmatrix}K(0)&K(t-s)\\ K(t-s)&K(0)\end{pmatrix}$,
@@ Appendix B Auxiliary Lemmas
Suppose that Assumptions B3, B6 and B8 hold. Then, it follows for the block diagonals $D_{n}:=\mathrm{diag}\left(h_{n}^{-1/2}I_{p},I_{q}\right)$ and $C_{n}:=\mathrm{diag}\left(nh_{n}I_{p},\sqrt{n}I_{q}\right)$ that

### v3

@@ Abstract
We propose a contrast-based estimation method for Gaussian processes with time-inhomogeneous drifts, observed under high-frequency sampling. The process is modeled as the sum of a deterministic drift function and a stationary Gaussian component with a parametric kernel. Our method constructs a local contrast function from adjacent increments, which avoids inversion of large covariance matrices and allows for efficient computation. We prove consistency and asymptotic normality of the resulting estimators under general ergodicity conditions. A distinctive feature of our approach is that the drift estimator attains a nonstandard convergence rate, stemming from the direct Riemann integrability of the drift density. This highlights a fundamental difference from standard estimation regimes. Furthermore, when the local contrast fails to identify all parameters in the covariance kernel, moment-based corrections can be incorporated to recover identifiability. The proposed framework is simple, flexible, and particularly well-suited for high-frequency inference with time-inhomogeneous structure.
@@ 1 Introduction
The asymptotic properties of the full maximum likelihood estimator (MLE) for Gaussian processes have been studied mainly in the spatial statistics literature (low-frequency sampling in the theoretical point of view). Bachoc [2] analyzed the role of spatial sampling under increasing- and fixed-domain asymptotics, establishing consistency and asymptotic normality under suitable identifiability conditions. More recently, Bachoc [3] provided a comprehensive introduction with proofs, emphasizing that, while increasing-domain asymptotics allow consistent estimation of all covariance parameters, under fixed-domain asymptotics only microergodic parameters can be consistently estimated.
However, these results are confined to spatial frameworks with increasing- or fixed-domain asymptotics. To the best of our knowledge, there is no corresponding asymptotic theory for the full MLE under high-frequency temporal sampling. The present paper fills this gap by establishing consistency and asymptotic normality of the MLE for Gaussian processes with time-dependent mean functions observed at high frequency.
@@ 2.1 Gaussian processes with time-inhomogeneous drifts
which is not smooth at $t=0$ since it includes $|t|$ in the exponent. However, we can approximate this kernel with a ‘mollifier’ such as
@@
$\lim_{n\to\infty}\left[\sum_{i=1}^{\infty}g(t_{i-1})h_{n}-\sum_{i=1}^{n}g(t_{i-1})h_{n}\right]=\lim_{n\to\infty}\int_{nh_{n}}^{\infty}g(s)\,\mathrm{d}s{\ =0}.$
@@ 3.1 Local-Gauss contrast
Suppose the same conditions as in Theorem 3.1. Suppose further B6 and B7. Then, the estimator $\widehat{\theta}_{n}$ is asymptotically normal under (2.2):
@@
$\Sigma(\theta_{0}):=\begin{pmatrix}\displaystyle-\partial_{t}^{2}K_{\sigma_{0}}(0)\left[2\int_{0}^{\infty}\{\partial_{\xi}\mu_{\xi_{0}}(t)\}^{\otimes 2}\,\mathrm{d}t\right]^{-1}&0\\ 0&V_{2}^{-1}(\sigma_{0})V_{1}(\sigma_{0})V_{2}^{-1}(\sigma_{0})\end{pmatrix}$.
@@ 3.2 Moment estimators
$\sqrt{n}\left(\widetilde{\sigma}_{n}-\sigma_{0}\right)\xrightarrow{d}\mathcal{N}\left(\mbox{$0$},\ A^{-1}\Gamma^{2}A^{-\top}\right),\quad n\to\infty.$
@@
  • $G(x,y)=(y-x){y^{3}}$ captures nonlinear interactions between local variation and the magnitude.
@@ 4.1 Drifted Gaussian processes with Gaussian Kernels
For separate estimation of $\alpha$ and $\beta$, we shall consider the method of moment. For example, we can use Lemmas B.5 and B.6 with $f(x,\theta)=x^{2}$ and $G(x,y,\theta)=(y-x)^{2}$, respectively: it follows for
@@
$\frac{1}{n}\sum_{i=1}^{n}(Y_{i-1}^{n})^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha;\quad{\frac{1}{nh_{n}^{2}}\sum_{i=1}^{n}(\Delta_{i}^{n}Y)^{2}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\alpha\beta};\quad$
@@ 4.3.2 Discussion of numerical experiments
Possible directions for improvement are as follows. (1) Random subsample averaging (Jackknife-after-bootstrap type): Generate random subsamples to estimate $\xi$, and use the average of these estimates. This can weaken the correlation between the estimation errors of $\xi$ and $(\alpha,\beta)$ without significantly increasing the variance of $\xi$, thus maintaining stability in finite samples more effectively than simple sample splitting. (2) Orthogonalized estimating functions using lagged covariances: Design the estimating functions for $\alpha$ and $\beta$ so that their gradient with respect to $\xi$ is zero (Neyman orthogonal), thereby ensuring that the estimation error of $\xi$ does not appear as a first-order term. In particular, use covariances at sufficiently large lags, rather than short differences, to extract information on $\alpha$ while attenuating the influence of the drift. (3) One-step stabilization: Starting from stable initial values of Case (I) type (consistent at rate $h_{n}^{-1/2}$ for $\xi$ and $n^{1/2}$ for $\alpha$ and $\beta$), apply a single Newton update to the joint estimating equations. This can correct the first-order dependence between $\xi$ and $(\alpha,\beta)$ and potentially mitigate the bias imbalance observed in finite samples. The goal here is not to improve efficiency, but to achieve better error propagation alignment and numerical stability.
@@ 6 Proofs of main theorems
$D_{n}(\widehat{\theta}_{n}-\theta_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}{\cal N}\left(\mbox{$0$},I^{-1}(\theta_{0})J(\theta_{0})I^{-1}(\theta_{0})\right)=\mathcal{N}(0,\Sigma(\theta_{0})).$
@@
$\Phi_{n}(\widetilde{\sigma}_{n})=\mbox{$0$}.$
@@
$\sqrt{n}\Phi_{n}(\sigma_{0})\stackrel{{\scriptstyle{\cal L}}}{{\longrightarrow}}\mathcal{N}(\mbox{$0$},\Gamma^{2}),\quad n\to\infty,$
@@
$\sqrt{n}\Phi_{n}(\sigma_{0})\xrightarrow{d}\mathcal{N}(\mbox{$0$},\Gamma^{2}).$
@@ Appendix A Limit theorems for stationary Gaussian processes
$(Z_{t},Z_{s})^{\top}\sim{\cal N}(\mbox{$0$},\Sigma(t,s)),\quad\Sigma(t,s)=\begin{pmatrix}K(0)&K(t-s)\\ K(t-s)&K(0)\end{pmatrix}$,
@@ Appendix B Auxiliary Lemmas
Suppose that Assumptions B3, B6 and B8 hold. Then, it follows for the block diagonals $D_{n}:=\mathrm{diag}\left(h_{n}^{-1/2}I_{p},{\sqrt{n}}I_{q}\right)$ and $C_{n}:=\mathrm{diag}\left(nh_{n}I_{p},{I_{q}}\right)$ that